OpenStudy (anonymous):

Which shows this equation in standard form? 6 – 2(x – y) = 3x + 5

OpenStudy (anonymous):

A. 5x – 2y = 1 B. 5x – 2y = –1 C. 5x + 2y = 1 D. 2x – 5y = 1

OpenStudy (anonymous):

@twinz1999 @imawesome101 help i only have a little time

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

heyyyyyyyyy

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

my twin lol :)

OpenStudy (anonymous):

creppy face :D plz help though

OpenStudy (anonymous):

hahahah ohh chucky!!!

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

cool u sure o u think

OpenStudy (anonymous):

I think im not sure yet

OpenStudy (anonymous):

i think C

OpenStudy (anonymous):

But im not positive

OpenStudy (anonymous):

ok thanks i have another will u help

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

ok, but im really bad.. im not even sure that right:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

wheres question lol XD

OpenStudy (anonymous):

oops lol :)

OpenStudy (anonymous):

lol wow xD

OpenStudy (anonymous):

hahahahaha wow chucky!!!:)

OpenStudy (anonymous):

What are the intercepts of this line? A. x-int: –0.5; y-int: –1 B. x-int: 0.5; y-int: 1 C. x-int: –0.5; y-int: 1 D. x-int: 0.5; y-int: –1

OpenStudy (anonymous):

i mean brad hehehe

OpenStudy (anonymous):

wheres the line??w

OpenStudy (anonymous):

im dumb

OpenStudy (anonymous):

no ur not, i do the same thing all the time :) hehehehe i think ur smart

OpenStudy (anonymous):

thanks and i think your beautiful

OpenStudy (anonymous):

all i think about is when is my race gonna start

OpenStudy (anonymous):

Awwwwwwww u made my day!!!!!!!!!!!!!!!! :)))

OpenStudy (anonymous):

ur sooo sweet

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

no no its so c

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

yeah ur right my dad just helped i have 2 more

OpenStudy (anonymous):

hahahahahaha @imawesome101

OpenStudy (anonymous):

Which line is a graph of the equation: –2x + 5y = 10? A. line a B. line b C. line c D. line d

OpenStudy (anonymous):

i don't know that.. we are learning the same thing tho lol :)

OpenStudy (anonymous):

o

OpenStudy (anonymous):

Which line is a graph of the equation x = 6 ? A. line a B. line b C. line c D. line d

OpenStudy (anonymous):

are those slops?

OpenStudy (anonymous):

u mean slopes and no its wait i got it

OpenStudy (anonymous):

yes, and ok coolios :)